# Quantifying the Impact of Incompleteness on Identifying and Interpreting Galaxy Protocluster Populations with the TNG-Cluster Simulation

### v1

@@ Abstract
We use the TNG-Cluster simulation to examine how stellar mass and star formation rate (SFR) incompleteness affect the identification of density peaks within galaxy protoclusters at different redshifts. We focus on 352 protoclusters, defined as the set of galaxies that will reside within the virialized region of a $z=0$ cluster halo with total mass $\sim 10^{14.3-15.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, and consider only galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ (our baseline) at any redshift. We find that ${M}_{\star}$-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR-limited ($\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) subpopulations only recover the true highest galaxy density peak in $\lesssim 40\%$ of cases within an accuracy of $1.0$ pMpc ($\sim 2-2.5\arcmin$) at $z>2$. We find that at $z>2$ the highest galaxy density peaks do not generally coincide with the highest dark matter or stellar mass density peaks, and this separation is larger than 0.5 pMpc in $\sim 80\%$ of cases. Consequently, the region surrounding the true highest galaxy density peaks are not generally sites of enhanced star formation or accelerated mass growth relative to the remainder of the protocluster. Lastly, we find that the $\sim 4\arcmin$ apertures typically used to define spectroscopically-confirmed protoclusters are generally much smaller than the $8\arcmin$ apertures needed to study the progenitors of the most massive galaxy clusters at $z>2$.
@@ 1 Introduction
While the late stages of cluster assembly have been extensively studied through observations of virialized (or near-virialized) clusters to $z\lesssim 1.5$ (e.g., Ebeling et al., 2001; Lubin et al., 2009; Wilson et al., 2009; Muzzin et al., 2012; Reichardt et al., 2013; Gonzalez et al., 2019; Golden-Marx et al., 2019; Balogh et al., 2021; Biviano et al., 2021), the pre-virialized or “protocluster” stage remains less well understood. This is partly because observations of virialized clusters provide limited insight into their early formation history. Transformative events such as mergers and dynamical relaxation significantly reshape clusters and their galaxy populations, often erasing evidence of their initial conditions — though dynamical analysis can still provide clues about their assembly history (Biviano & Katgert, 2004). This challenge is further compounded by the difficulty of identifying and characterizing galaxy protoclusters, which lack clear markers characteristic of virialized galaxy clusters at $z\lesssim 2$, such as an established ICM or a prominent red sequence. For example, the ICM is crucial for identifying and characterizing galaxy clusters, as it emits X-rays via thermal bremsstrahlung (free-free emission) and produces a distinctive decrement in the cosmic microwave background (CMB) at submillimeter wavelengths through the thermal Sunyaev-Zel’dovich (SZ) effect (Sunyaev & Zeldovich, 1970), caused by high-energy ICM electrons upscattering CMB photons. These signatures of ICM physics are essential for detecting galaxy clusters in X-ray (e.g., Takey et al., 2016; Koulouridis et al., 2021; Xu et al., 2022) and SZ (e.g., Bleem et al., 2015, 2024; Klein et al., 2024) surveys.
@@
The current sample of spectroscopically-confirmed galaxy protoclusters remains generally sparse and heterogeneously selected. While some consistent trends have emerged – such as protocluster environments, relative to coeval field populations, being regions of accelerated galaxy evolution (Steidel et al., 2005; Hatch et al., 2011; Koyama et al., 2013; Cooke et al., 2014; Shimakawa et al., 2018a; Forrest et al., 2024; Helton et al., 2024a), enhanced star formation rates (Dannerbauer et al., 2014; Hayashi et al., 2016; Miller et al., 2018), enhanced AGN activity (Digby-North et al., 2010; Lemaux et al., 2014; Tozzi et al., 2022; Shah et al., 2024), enhanced merger rates (Hine et al., 2016; Liu et al., 2023; Giddings et al., 2025), and sites of a reversal in the star formation rate–density relation (Tran et al., 2010; Popesso et al., 2011; Lemaux et al., 2022) – other trends remain contentious or poorly understood. For instance, it is uncertain whether protoclusters are metal-enriched, metal-deficient, or neither (Kulas et al., 2013; Alcorn et al., 2019; Sattari et al., 2021), and the extent to which protocluster environments suppress or “quench” star formation is still unclear (e.g., Kubo et al., 2021; Shi et al., 2021b; Alberts & Noble, 2022; McConachie et al., 2022; Tanaka et al., 2024; Edward et al., 2024; Espinoza Ortiz et al., 2024; Naufal et al., 2024; McConachie et al., 2025).
More broadly, our understanding of protoclusters remains largely phenomenological, with investigations primarily limited to characterizing differences between protoclusters and coeval field populations, while the physics driving this differential evolution remains elusive. This limitation arises not only from small and heterogeneous sample sizes, which hinder population-level analyses, but also from fundamental challenges in identifying protoclusters, constructing unbiased samples, and acquiring sufficient spectroscopic data over large areas to map their full extent and assemble a tracer population that is representative of the underlying galaxy distribution. However, efforts to build large, uniformly selected, and spectroscopically-confirmed protocluster samples are underway (Hung et al., 2025). These efforts, combined with ever-growing spectroscopic catalogs in legacy fields Khostovan et al. (2025) and ongoing and upcoming surveys with next-generation observatories – including JWST (Li et al., 2024), Euclid (Euclid Collaboration et al., 2025a), LSST (Ivezić et al., 2019; Gully et al., 2024), and Roman (Rudnick et al., 2023) – will play a significant role in advancing this field.
@@
Lim et al. (2021) used hydrodynamical simulations and empirical models to show that simulated protoclusters at $z>2$ underestimate observed star formation rates by at least a factor of three, attributing this to limitations in numerical resolution. However, Gouin et al. (2022), using the TNG300 simulation from the IllustrisTNG project (Nelson et al., 2019a), found that simulated protoclusters reproduce the star formation rates, stellar masses, and galaxy richness observed in Planck-selected high-redshift protoclusters. Lim et al. (2024) used the FLAMINGOS simulation suite (Schaye et al., 2023) to investigate how variations in aperture definitions affect mass estimates of protoclusters, finding that common observational aperture choices can introduce biases of up to an order of magnitude in total mass (baryonic and dark matter) estimates at $z\lesssim 4$. Using the DIANOGA zoom-in cosmological hydrodynamical simulations (Bonafede et al., 2011; Rasia et al., 2015), Esposito et al. (2025) found enhanced star formation suppression in protocluster galaxies at $z=2.2$, particularly in the most massive halos.
@@
Other studies have leveraged constrained and large-scale simulations to investigate analogs of observed protoclusters. Ata et al. (2022) used constrained cosmological simulations, designed to match the observed galaxy distribution in the COSMOS field (Scoville et al., 2007), to confirm that many of these systems will collapse into massive clusters by $z=0$. Remus et al. (2023) utilized the Magneticum Pathfinder simulations (Dolag et al., 2016) to analyze the evolutionary history of protoclusters similar to SPT2349-56 (Miller et al., 2018) at $z\sim 4$, concluding that properties such as virial mass, star formation rate (SFR), stellar mass, and galaxy richness do not strongly correlate with the final cluster mass at $z=0$, and that these protoclusters will not be among the top ten most massive clusters at $z=0.2$. Yajima et al. (2022) used the FOREVER22 simulation to study the formation of supermassive black holes (SMBHs) and bright SMGs in SSA22 protocluster analogs at $z\sim 3$, finding that SMBHs form in the most massive halos (${M}_{\rm{halo}}\sim 10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$) and grow rapidly until their stellar mass exceeds ${M}_{\star}\gtrsim 10^{10}\leavevmode\nobreak\ {\rm M}_{\odot}$, after which feedback suppresses further growth, while dusty starburst galaxies emerge in massive halos (${M}_{\rm{halo}}\gtrsim 10^{13}\leavevmode\nobreak\ {\rm M}_{\odot}$) in the protocluster core.
@@
In this work we use the TNG-Cluster simulation Nelson et al. 2024, which offers a unique combination of a large sample of very massive galaxy clusters (352 with $M_{\rm{200}}>10^{14.3}\leavevmode\nobreak\ {\rm M}_{\odot}$ at $z=0$) and high baryonic mass resolution ($1.2\times 10^{7}\leavevmode\nobreak\ {\rm M}_{\odot}$), to explore potential biases in observed protocluster populations. Specifically, we investigate how often the angular sizes used to probe spectroscopically-confirmed protoclusters align with their theoretically expected spatial extents. We also evaluate how stellar mass and SFR incompleteness affect the identification of the highest density peaks and compare the average properties of protocluster galaxies within and beyond these peaks.
@@
This paper is structured as follows. In §2 we describe the TNG-Cluster simulation, define our protocluster samples, and explain our methods for quantifying the radial extent of protoclusters and their density fields. In §3 we present a comparison of predicted and observed sizes of galaxy protoclusters, quantify the impact of stellar mass and SFR incompleteness on identifying the highest density peaks in protoclusters, and compare the properties of protocluster galaxies within and beyond the highest density peak. Finally, in §4 we discuss the implications of our results and provide considerations for observational studies of galaxy protoclusters. We summarize our findings in §5.
@@ 2.1 TNG-Cluster Simulation
Figure 2: The projected spatial distribution of six clusters from the TNG-Cluster simulation, selected to fall within the three halo mass bins defined in Fig. 1. Each row represents an individual cluster, while each column corresponds to a different redshift snapshot. Black circles denote galaxies from our baseline population – galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within the virial radius ($R_{\rm{200c}}$, red circle) by $z=0$ – whereas gray circles represent galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside between $R_{\rm{200c}}<R<10\leavevmode\nobreak\ R_{\rm{200c}}$ at $z=0$.
@@ 2.2 Simulated Cluster Sample
In this study we analyze all $352$ galaxy clusters from TNG-Cluster to investigate their assembly histories and the influence of local galaxy density in protocluster environments on the properties of their constituent galaxies. In the left panel of Fig. 1 we show the halo mass assembly histories of all clusters (gray lines). The solid colored lines indicate the median halo mass for clusters grouped by their halo mass at $z=0$ (i.e., $M_{\rm{200}}^{z=0}$), with the shaded bands representing the $68\%$ quantiles.
A striking feature of this sample is the diversity in cluster assembly histories – some clusters accumulate their mass rapidly at early times, while others grow more gradually, only exceeding $10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$ at later times. However, as shown in the right panel of Fig. 1, a clear trend emerges when comparing the redshift at which a cluster first reaches the characteristic cluster mass scale ($\geq 10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$) and its halo mass at $z=0$. This trend reflects the hierarchical nature of structure formation, where more massive halos at $z=0$ tend to form earlier (though this trend reverses if “formation” is instead defined using a fractional mass threshold; Wechsler et al., 2002; Nadler et al., 2023). It also motivates partitioning the protocluster sample based on $M_{\rm{200}}^{z=0}$ and defining $z\sim 2$ as the transition epoch when the first halos surpass the characteristic cluster halo mass scale, marking the transition from galaxy protoclusters to clusters. Notably, the most distant known cluster, CL J1001 (Wang et al., 2016, 2018), located at $z=2.51$, is consistent with the redshift at which the first clusters in the TNG-Cluster simulation cross the canonical halo mass threshold of $M_{\rm{200}}>10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$.

### v2

@@ Abstract
We use the TNG-Cluster simulation to investigate how stellar mass and star formation rate (SFR) incompleteness affect the identification of density peaks within galaxy protoclusters at different redshifts. Our analysis focuses on a sample of $352$ protoclusters, defined as the progenitor populations of galaxies that reside within the virialized region of $z=0$ clusters with $M_{\rm{200}}^{z=0}\sim 10^{14.2-15.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. For comparison, we define our “baseline” protocluster population as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ at any redshift. We find that ${M}_{\star}$-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR-limited ($\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) subpopulations recover the baseline highest galaxy density peak in roughly $\sim 60\%$ of cases within an accuracy of $1.0$ pMpc (corresponding to an angular scale of $\sim 2-2.5\arcmin$) at $z>2$. This recovery fraction drops to $\sim 40-50\%$ when restricting to galaxies with ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. We find that the baseline highest galaxy density peaks typically coincide with the highest dark matter and stellar mass density peaks, with separations less than $0.5$ pMpc in $\sim 60-75\%$ of cases at $z>2$. This agreement drops to $\sim 45-50\%$ when restricting to galaxies with ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. These results indicate that identifying the densest regions of protoclusters – i.e., the core – is highly sensitive to stellar mass and SFR completeness limits. Nevertheless, at $z>2$ we find that the baseline highest galaxy density peaks are generally sites of enhanced star formation and accelerated mass growth relative to the remainder of the protocluster, consistent with some observational studies.
@@ 1 Introduction
While the late stages of cluster assembly have been extensively studied through observations of virialized (or near-virialized) clusters to $z\lesssim 1.5$ (e.g., Ebeling et al., 2001; Lubin et al., 2009; Wilson et al., 2009; Muzzin et al., 2012; Reichardt et al., 2013; Gonzalez et al., 2019; Golden-Marx et al., 2019; Balogh et al., 2021; Biviano et al., 2021; Gully et al., 2025; Hewitt et al., 2025), the pre-virialized or “protocluster” stage remains less well understood. This is partly because observations of virialized clusters provide limited insight into their early formation history. Transformative events such as mergers and dynamical relaxation significantly reshape clusters and their galaxy populations, often erasing evidence of their initial conditions — though dynamical analysis can still provide clues about their assembly history (Biviano & Katgert, 2004). This challenge is further compounded by the difficulty of identifying and characterizing galaxy protoclusters, which lack clear markers characteristic of virialized galaxy clusters $z\lesssim 2$, such as an established ICM (see Di Mascolo et al., 2023; Lepore et al., 2024, for the detection of a “proto-ICM” in the Spiderweb protocluster) or a prominent red sequence. For example, the ICM is crucial for identifying and characterizing galaxy clusters, as it emits X-rays via thermal bremsstrahlung (free-free emission) and produces a distinctive decrement in the cosmic microwave background (CMB) at submillimeter wavelengths through the thermal Sunyaev-Zel’dovich (SZ) effect (Sunyaev & Zeldovich, 1970), caused by high-energy ICM electrons upscattering CMB photons. These signatures of ICM physics are essential for detecting galaxy clusters in X-ray (e.g., Takey et al., 2016; Koulouridis et al., 2021; Xu et al., 2022) and SZ (e.g., Bleem et al., 2015, 2024; Klein et al., 2024) surveys.
@@
The current sample of spectroscopically confirmed galaxy protoclusters remains generally sparse and heterogeneously selected. While some consistent trends have emerged – such as protocluster environments, relative to coeval field populations, being regions of accelerated galaxy evolution (Steidel et al., 2005; Hatch et al., 2011; Koyama et al., 2013; Cooke et al., 2014; Shimakawa et al., 2018; Forrest et al., 2024; Helton et al., 2024a; Gururajan et al., 2025), enhanced star formation rates (Dannerbauer et al., 2014; Hayashi et al., 2016; Miller et al., 2018), enhanced AGN activity (Digby-North et al., 2010; Lemaux et al., 2014; Tozzi et al., 2022; Shah et al., 2024), enhanced merger rates (Hine et al., 2016; Liu et al., 2023; Giddings et al., 2025), and sites of a reversal in the star formation rate–density relation (Tran et al., 2010; Popesso et al., 2011; Lemaux et al., 2022) – other trends remain contentious or poorly understood. For instance, it is uncertain whether protoclusters are metal-enriched, metal-deficient, or neither (Kulas et al., 2013; Alcorn et al., 2019; Sattari et al., 2021), and the extent to which these early dense environments suppress or “quench” star formation is still unclear (e.g., Kubo et al., 2021; Shi et al., 2021b; Alberts & Noble, 2022; McConachie et al., 2022; Tanaka et al., 2024; Edward et al., 2024; Espinoza Ortiz et al., 2024; Xie et al., 2024; Szpila et al., 2025; McConachie et al., 2025).
More broadly, our understanding of protoclusters remains largely phenomenological, with investigations primarily limited to characterizing differences between protoclusters and coeval field populations, while the physics driving this differential evolution remains elusive. This limitation arises not only from small and heterogeneous sample sizes, which hinder population-level analyses, but also from fundamental challenges in identifying protoclusters, constructing unbiased samples, and acquiring sufficient spectroscopic data over large areas to map their full extent and assemble a tracer population that is representative of the underlying galaxy distribution. However, efforts to build large, uniformly selected, and spectroscopically confirmed protocluster samples are underway (Hung et al., 2025). These efforts, combined with ever-growing spectroscopic catalogs in legacy fields Khostovan et al. (2025) and ongoing and upcoming surveys with next-generation observatories – including JWST (Li et al., 2025), Euclid (Euclid Collaboration et al., 2025a), LSST (Ivezić et al., 2019; Gully et al., 2024), and Roman (Rudnick et al., 2023) – will play a significant role in advancing this field.
@@
More recent hydrodynamical simulations have provided additional insights. Lim et al. (2021) used simulations and empirical models to show that theoretical models can underpredict star formation rates in protoclusters by a factor of ten, primarily due to limitations in numerical resolution. However, Gouin et al. (2022), using the TNG300 simulation from the IllustrisTNG project (Nelson et al., 2019a), found that simulated protoclusters reproduce the star formation rates, stellar masses, and galaxy richness observed in Planck-selected high-redshift protoclusters. Similarly, using The Manhattan Suite Rennehan (2024) was able to reproduce extremely star-bursting protoclusters similar to SPT2349-56 (Miller et al., 2018), which has an estimated total $\mathrm{SFR}$ of $\sim 10,000\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. Lim et al. (2024) used the FLAMINGOS simulation suite (Schaye et al., 2023) to investigate how variations in aperture definitions affect mass estimates of protoclusters, finding that common observational aperture choices can introduce biases of up to an order of magnitude in total mass (baryonic and dark matter) estimates at $z\lesssim 4$. Using the DIANOGA zoom-in cosmological hydrodynamical simulations (Bonafede et al., 2011; Rasia et al., 2015), Esposito et al. (2025) found enhanced star formation suppression in protocluster galaxies at $z=2.2$, particularly in the most massive halos.
@@
Other studies have leveraged constrained and large-scale simulations to investigate analogs to observed protoclusters. Ata et al. (2022) used constrained cosmological simulations, designed to match the observed galaxy distribution in the COSMOS field (Scoville et al., 2007), to confirm that many of these systems will collapse into massive clusters by $z=0$. Remus et al. (2023) utilized the Magneticum Pathfinder simulations (Dolag et al., 2016) to analyze the evolutionary history of protoclusters similar to SPT2349-56 (Miller et al., 2018) at $z\sim 4$, concluding that properties such as virial mass, star formation rate (SFR), stellar mass, and galaxy richness do not strongly correlate with the final cluster mass at $z=0$, and that these protoclusters will not be among the top ten most massive clusters at $z=0.2$. Yajima et al. (2022) used the FOREVER22 simulation to study the formation of supermassive black holes (SMBHs) and bright SMGs in SSA22 protocluster analogs at $z\sim 3$, finding that SMBHs form in the most massive halos (${M}_{\rm{halo}}\sim 10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$) and grow rapidly until their stellar mass exceeds ${M}_{\star}\gtrsim 10^{10}\leavevmode\nobreak\ {\rm M}_{\odot}$, after which feedback suppresses further growth, while dusty starburst galaxies emerge in massive halos (${M}_{\rm{halo}}\gtrsim 10^{13}\leavevmode\nobreak\ {\rm M}_{\odot}$) in the protocluster core.
@@
In this work we use the TNG-Cluster simulation Nelson et al. 2024, which offers a unique combination of a large sample of very massive galaxy clusters (352 with $M_{\rm{200}}>10^{14.2}\leavevmode\nobreak\ {\rm M}_{\odot}$ at $z=0$) and high baryonic mass resolution ($1.2\times 10^{7}\leavevmode\nobreak\ {\rm M}_{\odot}$), to explore potential biases in observed protocluster populations. Specifically, we investigate how often the angular sizes used to probe spectroscopically confirmed protoclusters align with their theoretically expected spatial extents. Moreover, since observed protoclusters are primarily probed via galaxy overdensities, incompleteness in stellar mass and/or SFR in the observed samples can impact our ability to identify and interpret these structures. We therefore evaluate how these sources of incompleteness affect the identification of the highest density peaks in simulated protoclusters and compare the average properties of protocluster galaxies within and beyond these peaks to determine if they are consistent with observed trends.
@@
This paper is structured as follows. In §2 we describe the TNG-Cluster simulation, define our protocluster samples, and explain our methods for quantifying the radial extent of protoclusters and their density fields. In §3 we present a comparison of predicted and observed sizes of galaxy protoclusters, quantify the impact of stellar mass and SFR incompleteness on identifying the highest density peaks in protoclusters, and compare the properties of protocluster galaxies within and beyond the highest density peak. Finally, in §4 we discuss the impact of galaxy selection on protocluster peak identification and offer suggestions for interpreting existing and future protocluster samples. We summarize our findings in §5.
@@ 2.1 TNG-Cluster Simulation
Figure 2: The projected spatial distribution of six clusters from the TNG-Cluster simulation, selected to fall within the three halo mass bins defined in Fig. 1. Each row represents an individual cluster, while each column corresponds to a different redshift snapshot. Black circles denote galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that are the progenitors of the galaxies that reside within the cluster’s virial radius ($R_{\rm{200c}}$, red circle) at $z=0$ – this defines our baseline population. Gray circles represent galaxies above the same stellar mass threshold that are the progenitors of the galaxies that reside between $R_{\rm{200c}}<R<10\leavevmode\nobreak\ R_{\rm{200c}}$ at $z=0$.
@@ 2.2 Simulated Cluster Sample
In this study we analyze all $352$ galaxy clusters from the TNG-Cluster simulation to investigate their assembly histories and the influence of local galaxy density in protocluster environments on the properties of their constituent galaxies. In the left panel of Fig. 1 we show the halo mass assembly histories of all clusters (gray lines). The solid colored lines indicate the median halo mass for clusters grouped by their halo mass at $z=0$ (i.e., $M_{\rm{200}}^{z=0}$), with the shaded bands representing the $68\%$ quantiles.
[…]
Rather than comparing progenitor masses at a fixed redshift, we can isolate the redshift at which a halo first reaches the characteristic cluster mass scale – i.e., $M_{\rm{200}}\geq 10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$. We show this in the right panel of Fig. 1, where we find a positive correlation between this formation redshift and halo mass at $z=0$, though the relationship exhibits substantial scatter. This trend reflects the hierarchical nature of structure formation, where the most massive halos at $z=0$ exhibit a tendency to begin forming earlier (though this trend reverses if “formation” is instead defined using a fractional mass threshold; Wechsler et al., 2002; Nadler et al., 2023). However, the strength of this correlation is likely enhanced by the TNG-Cluster selection function, which includes only clusters with $M_{\rm{200}}^{z=0}>10^{14.2}\leavevmode\nobreak\ {\rm M}_{\odot}$.
Nevertheless, this trend motivates partitioning the protocluster sample based on $M_{\rm{200}}^{z=0}$ and defining $z\sim 2$ as the transition epoch when the first halos surpass the characteristic cluster halo mass scale, marking the transition from galaxy protoclusters to clusters. Notably, the most distant known cluster, CL J1001 (Wang et al., 2016, 2018), located at $z=2.51$, is consistent with the redshift at which the first clusters in the TNG-Cluster simulation cross the canonical halo mass threshold of $M_{\rm{200}}>10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$.